# Extracting scattering phase shift in quantum mechanics on quantum computers

### v1

@@ I Introduction
Computationally, all of the approaches are so demanding that lattice QCD simulations have been constantly pushing the limits of high-performance computing. With the advent of quantum computing, it is natural for the field to explore its potential to speed up the simulations. In addition to larger lattices, near or at physical pion masses, and multi-baryon systems, quantum computing offers new prospects to overcome intrinsic limitations facing classical simulations, such as no access to real-time dynamics, and the sign problem in finite-density systems [27, 28]. Various ideas have been proposed in nuclear physics and lattice QCD in general [29, 30, 31, 32]; and scattering in particular [33, 34, 35, 36, 37, 38, 39, 40, 41].
@@ II Scattering in 1D Quantum Mechanical model
By discretizing $x$ into $x_{\alpha}=-\frac{L}{2}+a\alpha$, where $a=\frac{L}{N-1}$ is lattice spacing and $\alpha\in[0,1,\cdots,N-1]$, and also replacing continuous derivative by
@@ II.1 Relating trapped dynamics to infinite volume scattering phase shift
Figure 1: Convergence of Eq.(19) in Euclidean time with $L=4$ (dashed red), 1010 (dashed red) and $L=\infty$ (solid black). Both repulsive potential $V_{0}=2$ (upper panel) and attractive potential $V_{0}=-0.5$ (lower panel) are demonstrated for parameters $m=1$ and $N=400$. The corresponding lattice spacing is $a=\frac{L}{N-1}=0.01,\,0.025$.
@@ II.2 Post data analysis
One idea is to smooth out the fast oscillatory behavior by averaging out over a short period of time that is larger than oscillation period, as was suggested in Ref. [26]. In the present work, we explore other possibilities of post-data analysis. Using the integral transform $i\int_{0}^{\infty}dte^{iEt}\left[\cdots\right]$, Eq.(11) can be brought into a compact form amenable to numerical verification,
@@ IV Numerics
Figure 13: Single-qubit demo plots for real (upper panel) and imaginary (lower panel) parts of $\triangle C(t)=C(t)-C_{0}(t)$ on IBM QPUs (black dots) vs. exact results (solid red and solid purple curves) for a repulsive potential $V_{0}=2$ and other parameters $L=4$, $m=1$, $\delta t=0.04$. The number of quantum measurements (shots number) is 1000.
@@
Figure 15: Two-qubit demo plots of real (upper panel) and imaginary (lower panel) parts of $\langle\alpha|e^{-i\hat{H}t}|\alpha\rangle$ with $\alpha=0$ on IBM QPUs without error mitigation (black dots) and with error mitigation (orange dots) vs. exact results (solid red and solid purple curves) vs. simulator results (blue dots) for a repulsive potential $V_{0}=2$ and parameters $L=4$, $m=1$, $\delta t=0.04$. The shots number is 1000.
@@
For two qubits, the quantum circuit of the lowest order Trotterization approximation of $U(\delta t)=e^{-i\hat{H}\delta t}$ is shown in Fig. 14. The demo plots of a single term $\langle\alpha|e^{-i\hat{H}t}|\alpha\rangle$ (we picked $\alpha=0$ in Eq.(45)) on IBM devices vs. exact results vs. simulated results are shown in Fig.15. We see that the results on Qiskit simulator without errors are consistent with the exact result, but the two-qubit results (black dots) on IBM devices failed completely, even at the lowest order of Trotterization and for a single term in Eq.(45). These measured expectation values collapse rapidly toward zero, losing any of the theoretical behavior, becoming indistinguishable from random noise after only a few Trotter steps. We utilized the mthree package [59] for error mitigation, shown as orange dots in Fig.15; no improvement is observed.
@@
In contrast, the two–qubit gate errors begin to show the sharp decoherence patterns that we see when running the demos on hardware. Due to the number of gate operations required by this lowest–order Trotterization, even the most optimistic two–qubit error patterns realistically achievable on current hardware simulate to strongly kill off any coherent dynamics needed for evolution. Because of the multiple entangling gates required by each Trotter step, the total two–qubit gate count grows faster than coherent phase information can be preserved as evolution time increases causing stochastic errors to become substantially more likely. The simulation, Fig. 18, implies that at least a part of this sharp drop to near random outcomes is coming from this fight between the required relatively long circuit depth compared and the currently achievable coherence times.
@@
We see this behavior as well when considering thermal relaxation errors in Fig. 19 throughout many different ranges of realistic gate lengths and relaxation times.
@@ Acknowledgements.
This research is supported by the U.S. National Science Foundation under grant PHY-2418937 (P.G.) and in part Grant No. PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) (P.G.), and the U.S. Department of Energy under grant DE-FG02-95ER40907 (F.L.). The work of YZ is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-AC02-06CH11357.
@@ Appendix C Extension to scalar field theory models
The discretized positions are defined by $x_{j}=aj$, where $j\in[0,N_{x}-1]$. The $a=\frac{L}{N_{x}-1}$ is lattice spacing, and $L$ is the size of periodic box along spatial extent. The $m$ and $\lambda$ refer to the mass and coupling strengths of scalar field $\hat{\phi}$ respectively. The conjugate momenta operator $\hat{\Pi}(x_{j})$ is given by

### v2

# Scattering phase shift in quantum mechanics on quantum computers

@@ I Introduction
Computationally, all of the approaches are so demanding that lattice QCD simulations have been constantly pushing the limits of high-performance computing. With the advent of quantum computing, it is natural for the field to explore its potential to speed up the simulations. In addition to larger lattices, near or at physical pion masses, and multi-baryon systems, quantum computing offers new prospects to overcome intrinsic limitations facing classical simulations, such as no access to real-time dynamics, and the sign problem in finite-density systems [27, 28]. Various ideas have been proposed in nuclear physics and lattice QCD in general [29, 30, 31, 32]; and scattering in particular [33, 34, 35, 36, 37, 38, 39, 40, 41]. Additionally, there have been rapid developments recently on the subject of reaction processes in quantum simulations, including (1) combined Variational Quantum Eigensolver and Lüscher- or BERW-like formula approach in Ref. [42, 43]; (2) measuring phase shift by wave packet time delay [44]; (3) reconstructing scattering amplitudes through the coupling of the particle with an ancillary spin-1/2 [45]; (4) radiative processes [46]; (5) inclusive process; and other ideas, e.g. [47, 48, 49, 50].
@@ II Scattering in 1D Quantum Mechanical model
By discretizing $x$ into $x_{\alpha}=-\frac{L}{2}+a\alpha$, where $a=\frac{L}{N}$ is lattice spacing and $\alpha\in[0,1,\cdots,N-1]$, and also replacing continuous derivative by
@@ II.1 Relating trapped dynamics to infinite volume scattering phase shift
Figure 1: Convergence of Eq.(19) in Euclidean time with $L=4$ (dashed red), 1010 (dashed red) and $L=\infty$ (solid black). Both repulsive potential $V_{0}=2$ (upper panel) and attractive potential $V_{0}=-0.5$ (lower panel) are demonstrated for parameters $m=1$ and $N=400$. The corresponding lattice spacing is $a=\frac{L}{N}=0.01,\,0.025$.
@@ II.2 Post data analysis
One idea is to smooth out the fast oscillatory behavior by averaging out over a short period of time that is larger than oscillation period, as was suggested in Ref. [26] (also see Ref.[60]). In the present work, we explore other possibilities of post-data analysis. Using the integral transform $i\int_{0}^{\infty}dte^{iEt}\left[\cdots\right]$, Eq.(11) can be brought into a compact form amenable to numerical verification,
@@ IV Numerics
Figure 13: Single-qubit demo plots for real (upper panel) and imaginary (lower panel) parts of $\triangle C(t)=C(t)-C_{0}(t)$ on IBM QPUs (black dots) vs. exact results (solid red and solid purple curves) for a repulsive potential $V_{0}=2$ and other parameters $L=8$, $m=1$, $\delta t=0.04$. The number of quantum measurements (shots number) is 1000.
@@
Figure 15: Two-qubit demo plots of real (upper panel) and imaginary (lower panel) parts of $\langle\alpha|e^{-i\hat{H}t}|\alpha\rangle$ with $\alpha=0$ on IBM QPUs without error mitigation (black dots) and with error mitigation (orange dots) vs. exact results (solid red and solid purple curves) vs. simulator results (blue dots) for a repulsive potential $V_{0}=2$ and parameters $L=5.33$, $m=1$, $\delta t=0.04$. The shots number is 1000.
@@
For two qubits, the quantum circuit of the lowest order Trotterization approximation of $U(\delta t)=e^{-i\hat{H}\delta t}$ is shown in Fig. 14. The demo plots of a single term $\langle\alpha|e^{-i\hat{H}t}|\alpha\rangle$ (we picked $\alpha=0$ in Eq.(45)) on IBM devices vs. exact results vs. simulated results are shown in Fig.15. We see that the results on Qiskit simulator without errors are consistent with the exact result, but the two-qubit results (black dots) on IBM devices failed completely, even at the lowest order of Trotterization and for a single term in Eq.(45). These measured expectation values collapse rapidly toward zero, losing any of the theoretical behavior, becoming indistinguishable from random noise after only a few Trotter steps. We utilized the mthree package [69] for error mitigation (also see a similar readout error mitigation approach in [70]), shown as orange dots in Fig.15; no improvement is observed.
@@
In contrast, the two–qubit gate errors begin to show the sharp decoherence patterns that we see when running the demos on hardware. Due to the number of gate operations required by this lowest–order Trotterization, even the most optimistic two–qubit error patterns realistically achievable on current hardware simulate to strongly kill off any coherent dynamics needed for evolution. Because of the multiple entangling gates required by each Trotter step, the total two–qubit gate count grows faster than coherent phase information can be preserved as evolution time increases causing stochastic errors to become substantially more likely. The simulation, Fig. 18, implies that at least a part of this sharp drop to near random outcomes is coming from this fight between the required relatively long circuit depth compared and the currently achievable coherence times. Readout errors up to 5% and single qubit gate errors at typical hardware rates (around 0.01% – 0.1%) produce only minor distortions, while the two–qubit gate errors at realistic values (around 0.25% – 1%) rapidly suppress the coherent evolution once several Trotter steps are applied. These results indicate that, for circuits of the depth considered here, the two–qubit gate errors would need to be reduced to roughly the level of $0.01\%$ or below for the signal to remain visible within a reasonable deviation at moderate evolution times. Our noise simulations allow us to estimate the approximate error levels required to preserve a visible signal in the correlator.
@@
We see this behavior as well when considering thermal relaxation errors in Fig. 19 throughout many different ranges of realistic gate lengths and relaxation times. The thermal noise simulations similarly indicate that realistic coherence times place strong limits on the usable circuit depth. Even for relatively optimistic parameters $(T1,T2)=(250\mu s,150\mu s)$, the signal decays rapidly once the two–qubit gate durations reach the 100–500ns range typical of current devices, indicating that substantially longer coherence times or shorter entangling gates would be required to preserve coherent evolution over multiple Trotter steps, on the order of 50ns for the two–qubit gates.
@@ Acknowledgements.
This research is supported by the U.S. National Science Foundation under grant PHY-2418937 (P.G.) and PHY-2531653 (P.G.) and in part Grant No. PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) (P.G.), and the U.S. Department of Energy under grant DE-FG02-95ER40907 (F.L.). The work of YZ is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-AC02-06CH11357.
@@ Appendix C Extension to scalar field theory models
The discretized positions are defined by $x_{j}=aj$, where $j\in[0,N_{x}-1]$. The $a=\frac{L}{N_{x}}$ is lattice spacing, and $L$ is the size of periodic box along spatial extent. The $m$ and $\lambda$ refer to the mass and coupling strengths of scalar field $\hat{\phi}$ respectively. The conjugate momenta operator $\hat{\Pi}(x_{j})$ is given by